Required information [The following information applies to the questions displayed below.] On October 29, Lobo Co. began oper1. Prepare journal entries to record above transactions and adjustments. View transaction list Journal entry worksheet 2 3 48 Record the estimated warranty expense at 8% of December sales. 9 Record the sales revenue of 140 razors for $9,800 cash. 102. How much warranty expense is reported for November and December? Warranty expense for November Warranty expense for Decemb3. How much warranty expense is reported for January? Warranty expense4. What is the balance of the Estimated Warranty Liability account as of December 31? Estimated warranty liability balance5. What is the balance of the Estimated Warranty Liability account as of January 31? Estimated warranty liability balance
Answer #1
1.1
Date General Journal Debit Credit
11-Nov Cash 4,900
Sales 4,900
11-Nov Cost of goods sold 1050 =70*15
Merchandise inventory 1,050
30-Nov Warranty expense 392 =4900*8%
Estimated warranty liability 392
9-Dec Estimated warranty liability 210 =14*15
Merchandise inventory 210
16-Dec Cash 14,700
Sales 14,700
16-Dec Cost of goods sold 3,150 =210*15
Merchandise inventory 3,150
29-Dec Estimated warranty liability 420 =28*15
Merchandise inventory 420
31-Dec Warranty expense 1,176 =14700*8%
Estimated warranty liability 1,176
1.2
Date General Journal Debit Credit
5-Jan Cash 9,800
Sales 9,800
5-Jan Cost of goods sold 2,100 =140*15
Merchandise inventory 2,100
17-Jan Estimated warranty liability 495 =33*15
Merchandise inventory 495
31-Jan Warranty expense 784 =9800*8%
Estimated warranty liability 784
2
Warranty expense for November 392 =4900*8%
Warranty expense for December 1176 =14700*8%
3 Warranty expense 784 =9800*8%
4
Warranty expense for November 392
Warranty expense for December 1176
Less: Cost of replacements -630 =(14+28)*15
Estimated Warranty Liability balance 938
5 Estimated Warranty Liability balance, beginning 938
Warranty expense for January 784
Less: Cost of replacements -495 =33*15
Estimated Warranty Liability balance 1227